OpenStudy (anonymous):

Use the properties of exponents to simplify the expression. Write with positive exponents. Assume that all variables represent positive real numbers. b^(3/7) times b^(8/7)

OpenStudy (anonymous):

\[\frac{3}{7}+\frac{8}{7}=\frac{3+8}{7}=\frac{11}{7}\]

OpenStudy (saifoo.khan):

\[\Large b^{\frac37 + \frac87}\]

OpenStudy (anonymous):

so your answer is \[b^{\frac{11}{7}}\]